One saw and 4 hammers cost $72. Two saws and 6 hammmers cost $114. How much does each item cost?
Answer by mananth(16946)

saws x $
hammers y $

1 x + 4 y = 72 .............1
2 x + 6 y = 114 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 2
-3 x -12 y = -216
4 x + 12 y = 228
Add the two equations
1 x = 12.00
/ 1
x = 12
plug value of x in (1)
1 x + 4 y = 72
12 + 4 y = 72
4 y = 72 -12
4 y = 60
y = 15
saws $ 12
hammers $ 15
m.ananth@hotmail.ca